When Grace had lunch at Subway, she decided to try and figure out how many possible sandwiches are possible. If she knows that a person could have 5 different types of bread in 3 sizes, 11 kinds of meat, 4 kinds of cheese, 12 types of veggies, with 7 types of spread, how many different sandwiches are available if a person chooses one of each of the above options?
Answer #1

SOLUTION:

From given data,

When Grace had lunch at Subway, she decided to try and figure out how many possible sandwiches are possible. If she knows that a person could have 5 different types of bread in 3 sizes, 11 kinds of meat, 4 kinds of cheese, 12 types of veggies, with 7 types of spread,

We need to find ,

how many different sandwiches are available if a person chooses one of each of the above options

where,

5 different types of bread

bread in 3 sizes

11 kinds of meat

4 kinds of cheese

12 types of veggies

7 types of spread

Number of sandwiches possible with one each of the above options

= 5 \times 3 \times 11 \times 4  \times 12 \times 7

= 55440

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.